OpenStudy (anonymous):

i don't understand the example:

OpenStudy (anonymous):

\[\left(\begin{matrix}\frac{ p+1 }{ 2 } \\ \end{matrix}\right)^{2}+(p+1)^{3}\] = \[\left(\begin{matrix}\frac{ p+1 }{ 2 } \\ \end{matrix}\right)^{2}.1 + \left(\begin{matrix}\frac{ p+1 }{ 2 } \\ \end{matrix}\right)^{2}\frac{ (p+1)^3 }{ \left(\begin{matrix}\frac{ p+1 }{ 2 } \\ \end{matrix}\right)^{2}}\]

OpenStudy (anonymous):

@cwrw238 @ganeshie8 @satellite73

OpenStudy (anonymous):

@Carl_Pham @experimentX @Sam.

OpenStudy (experimentx):

huh ... that's pretty obvious \[ a + b = a+ {a \times b \over a}\]

OpenStudy (experimentx):

let \[ \left( p+1 \over 2\right)^2 = a \text{ and } (p+1)^3 = b \]

OpenStudy (anonymous):

but why are they dividing by a...

OpenStudy (anonymous):

wait @experimentX , so they divide by a so that they can have a common denominator?

OpenStudy (experimentx):

really i don't know ... probably so that they can take common well i thought i would easily copy your latex ... but you seem to type complex expression.

OpenStudy (experimentx):

\[ \left( p+1 \over 2\right)^2 \left( 1 + {(p+1)^3 \over \left( p+1 \over 2\right)^2}\right)\]

OpenStudy (anonymous):

its fine , thank you for the help :)